# PAT: Pixel-wise Adaptive Training for Long-tailed Segmentation

### v2

@@ 5 Experimental Evaluations and Discussion
To evaluate the performance among methods, we conducted various experiments on three popular datasets including OxfordPet [19], CityScapes [5], and NyU [16] whose frequency of classes is considerably imbalanced in each sample with even missing classes, which refers to Section 5.1. The training, validating, and testing ratios are 0.8, 0.1, and 0.1, respectively, applied to all considered datasets. In this paper we compare the proposed method with Focal [13], Class Balance Loss (CB) [6], the combination of CB and Focal (CBFocal) [6], Balance Meta Softmax (BMS) [21], Label Distribution Aware Margin Loss (LDAM) [1], and Balance Logits Variation (BLV) [26] based three main metrics: mean Intersection over Union (mIoU %), pix acc (Pix Acc %), and Dice Error (Dice Err), which are the three main metrics for evaluating performance of segmentation model. Otherwise, all experiments are trained with a fixed seeding number of 0. The number of rounds of all experiments is fixed to 100 rounds.
To conduct a fair comparison among methods, we conduct the overall training procedure by using SegNet[2] model architecture and tuning the hyperparameter of each loss function to find the best case. We also experiment with vanilla Cross-entropy loss function (CE). Specifically, 1) Focal and the combination of Class Balance and Focal are trained with different values of $\gamma\in\{0.5,1,2,3,4,5\}$ [13, 6]. 2) In the LDAM loss function, we set the parameter $\mu$ of $0.5$ as the default setting in [4] and trained with different scale $s\in\{10,20,30,50\}$. 3) For the BLV method, we applied multiple types of distribution including Gaussian, Uniform, and Xaviver along with different standard deviation $\sigma\in\{0.5,1,2\}$. 4) Hyper-parameter tuning is conducted on our proposed method with various values of $T\in\{5,10,20,50\}$ (refer to Section 5.2.2).
5.1 Comparisons to state-of-the-arts (SOTA)
In Table 1, the PAT outperforms the others in almost all settings from $0.09\%$ to $2.2\%$ in mIoU and from $0.07\%$ to $0.36\%$ in pix acc, respectively. While the dice error of BLV and PAT are identically 0.23 in the Oxford dataset, PAT’s dice error in CityScapes and NyU datasets (two higher number of categories dataset) decreased around 0.02 and 0.39, respectively. Regarding Fig. 4, models trained by PAT can segment objects with mIoU of 76.22% and pix acc of 85.80%, increasing 2.2% and 0.36%, respectively. In detail, these figures illustrate visually how PAT can balance the class-wise consideration.
First, PAT can enhance models on detecting long-tailed rare objects by putting a higher weight on the class-wise loss function (refer to Figs. 2(a), 2(c)). In the second column of Fig. 4 (Munich domain), under a long-tailed scenario where most parts of the image are road and sky, while the building owns just a small proportion, the model trained by PAT can segment fully the road and sky as well as most of the building. This phenomenon is also true with models trained on OxfordPetIII (refer to Fig. 5).
Second, as aforementioned, the proposed method can tackle underlying issues in detecting long-tailed rare objects by using scalarization coefficients without forgetting the well-classified categories (refer to Fig. 2(c)). Visualization performance in Berlin and Leverkusen (refer to Fig. 4), are two typical examples. Other SOTAs tend to misclassify the sky and the road, even though the sky and road account for a considerable proportion. This phenomenon is not an exception, which also appears in the Leverkusen example, where other SOTAs misclassify car masks, which take $40\%$ in mask size of the whole portion.
Table 1: Overall Performance of 8 baselines (i.e. Vanilla Softmax, Focal, Class Balance Loss, Class Balance Focal) and proposed method among three different scenarios including OxfordPetIII, CityScape, and NyU datasets. In detail, the bold number indicates the best performance, while $\uparrow$ and $\downarrow$ show "higher is better" and "lower is better", respectively. Note that all experiments shown in this table are trained using SegNet[2] architecture.
@@ 5.1 Comparisons to state-of-the-arts (SOTA)
CE
76.14 91.21 0.23 73.83 85.31 0.66 18.05 53.50 1.80
Focal ($\gamma=2$)
75.76 91.17 0.30 74.02 85.44 0.56 15.14 51.07 1.75
CB
76.60 90.90 0.20 72.26 81.33 0.69 18.56 52.17 1.94
CBFocal ($\gamma=2$)
76.02 90.54 0.26 71.17 80.70 0.72 17.31 50.46 1.76
BMS
13.22 25.45 1.28 8.15 11.80 3.08 12.27 22.84 2.40
LDAM ($\mu=0.5$, $s=20$)
75.43 90.97 0.78 74.80 85.20 2.27 19.59 52.59 2.30
BLV (Gaussian, $\sigma=0.5$)
76.24 91.22 0.23 74.21 85.37 0.53 18.37 52.62 1.90
@@
76.22
$\uparrow$ 2.2
85.80
$\uparrow$ 0.36
0.51
$\downarrow$ 0.02
21.41
$\uparrow$ 2.85
55.57
$\uparrow$ 2.07
1.36
$\downarrow$ 0.39
@@
Figure 4: Segmentation visualization of models trained by the proposed method PAT and other baselines on the CityScapes dataset. The two first rows are the original image on the test set and its corresponding ground truth, respectively. The third row illustrates the performance in visualization of the proposed method. We sample four typical examples from four corresponding different domains containing Berlin, Munich, Bielefeld, and Leverkusen.
5.1.1 OxfordPetIII Dataset [19].
Oxford dataset contains 37 categories of dogs and cats with roughly 200 images for each type. The mask ground truth of each image includes three classes: background, boundary, and main body of the animals. All images are collected with a high resolution of $640\times 340$ pixels, which is then resized to $256\times 256$ pixels. In this dataset, the biggest obstacle is to segment the boundary of the animal which is very small and not easily distinguishable. In training the image is preprocessed by diving the max value of that image. The labels, on the other hand, are converted into tensors of one-hot vectors at pixel levels.
5.1.2 CityScapes Dataset [5].
CityScapes dataset contains roughly 1500 images with a high resolution of $1024\times 512$ pixels, which is then resized to $512\times 256$ pixels. Note that adjusting the image size does not change the proportions among classes. The ground truth in this dataset is labeled in fine mode, in which there are no boundaries among classes. This dataset faced a big issue in the imbalance among classes (refers to Fig. 2). There are potentially many categories that only take a small account of the mask as well as a large number of ones that are not included.
5.1.3 NyU Dataset [16].
The NyU dataset contains roughly 1000 samples, with 14 different categories. Each sample contains an image and label whose size of $340\times 256$ pixels. As in the CityScapes dataset, NyU also faces a big challenge in segmenting long-tailed rare objects. In the preprocessing stage, the image is normalized into the range of $[0,1]$, and its corresponding ground truth is converted into a tensor of one-hot vectors at pixel levels.
@@ 5.2 Ablation studies
To guarantee the proposed method’s adaptability and independence in different deep architecture designs[1, 18, 29, 24, 12, 27], we analyze the performance on 3 additional different model structures containing UNet[22], Attention UNet[17] (AttUNet), and Nested UNet[33] (UNet++). The hyperparameter settings are presented in Table 1. The quantitative performance results of this ablation test are shown in Table 2. In detail, we compare the performance among methods conducted in each model. The corresponding colors of UNet, AttUNet, and UNet++ are blue, green, and yellow, respectively, which indicate the best cases.
According to Table 2, PAT outperforms the other baselines in CityScapes experiments, which is the highest number of classes dataset. Especially in Unet++ experiments, the mIoU and the pix acc of CityScapes are above $75\%$ and $85\%$, compared to under the two aforementioned values in BLV[26] and LDAM[4]. Although the Class-Balance loss function tends to work well with the OxfordPetIII dataset (3 distinct categories), which reaches $78.71\%$ and $76.60\%$ of mIoU (refer to Tables 12) trained by UNet and SegNet, respectively, CB exhibit low performance on datasets with higher number of categories.
LDAM and BLV, otherwise, achieve higher performance as opposed to PAT from $0.01\%$ to $0.04\%$ in the NyU dataset. In detail, LDAM achieves $55.22\%$ of pix acc using AttUNet and BLV reaches $54.67\%$ and $54.93\%$ of pix acc using UNet and UNet++, respectively. However, performance in mIoU of PAT outperforms LDAM and BLV from $1\%$ to $3\%$ (refer to Table 2). mIoU is more crucial than pix acc, which not only indicates the correctness but also the completeness of segmented objects [28]. PAT, otherwise, outperforms other baselines in both the CityScapes dataset and the NyU dataset on various types of model architecture, which achieve $74.85\%$, $74.57\%$, and $75.24\%$ of mIoU and $21.18\%$, $21.44\%$, and $20.66\%$ of mIoU on UNet, AttUNet, and UNet++, respectively. Figure 5 illustrates that PAT can segment the object completely, compared to other baselines that tend to misclassify the objects’ boundary.
Table 2: Quantitative comparisons between baselines and the proposed method in three datasets and three model architecture designs: UNet, Attention UNet, and Nested UNet. Note that the blue, green, and yellow colors indicate the outperforming cases trained on UNet, Attention UNet, and Nested UNet, respectively.
OxfordPetIII CityScapes NyU
@@
UNet 78.61 92.40 74.08 83.63 15.36 50.97
AttUnet 78.83 92.52 73.28 87.33 16.46 51.35
CE UNet++ 78.18 92.30 72.91 82.87 16.48 51.42
UNet 78.27 92.24 73.53 83.69 15.42 51.03
AttUnet 78.31 92.30 73.93 83.10 16.46 51.32
Focal ($\gamma=2$) UNet++ 77.78 92.15 74.21 83.52 16.72 51.41
UNet 78.71 91.99 73.58 83.19 20.18 54.91
AttUnet 79.26 92.31 74.47 83.76 19.07 54.66
CB UNet++ 78.51 91.96 74.85 83.66 19.17 53.23
UNet 78.13 91.61 73.74 83.01 17.82 54.14
AttUnet 78.82 92.02 73.42 83.18 18.76 54.26
CBFocal ($\gamma=2$) UNet++ 77.76 91.54 72.58 82.45 17.09 53.02
UNet 78.68 92.44 73.42 83.81 18.04 54.33
AttUnet 78.71 92.51 73.73 83.46 19.81 55.22
LDAM ($\mu=0.5$, $s=20$) UNet++ 78.06 92.20 73.55 84.64 19.04 52.99
UNet 78.25 92.25 74.72 84.74 18.32 54.67
AttUnet 78.97 92.46 74.45 84.56 19.45 54.72
BLV (Gaussian, $\sigma=0.5$) UNet++ 78.37 92.23 74.93 84.86 19.81 54.93
UNet 78.63 92.27 74.85 85.51 21.18 54.22
AttUnet 79.37 92.72 74.57 85.56 21.44 54.41
PAT (Ours) ($T=20$) UNet++ 79.14 92.51 75.24 85.80 20.66 54.86
@@
In this part, we perform various experiments of PAT with different values of temperature $T$, which are $\{5,10,20,50\}$ (refers to Table 3). This ablation test analyzes how temperature parameter $T$ affects the performance of the segmentation model. To make a fair comparison, we conduct all experiments with three related datasets as mentioned in Section 5.
In OxfordPetIII-related experiments, adjusting the temperature parameter $T$ does not change the performance of the model which the performance is ranging under 77% and roughly 78.5%, using SegNet and UNet, respectively. On the other hand, the CityScapes dataset whose number of classes is much higher, which is 20, compared to 3 in OxfordPetIII, is indicated to be more sensitive to the temperature parameter $T$. According to Table 3, The mIoU and pix acc of CityScapes increase from $73.86\%$ to $76.22\%$ and from $84.82\%$ to $85.76\%$, respectively using SegNet model architecture. SegNet model performance in the NyU dataset, on the other hand, achieves from $19.35\%$ to $21.41\%$ in mIoU and from $53.39\%$ to $54.29\%$ in pix acc. The UNet model is not an exception, whose performance on CityScapes and NyU increases $0.7\%$ in mIoU, $1.7\%$ in pix acc, and $3.1\%$ in mIoU, $2.2\%$ in pix acc, respectively. These quantitative results suggest that the temperature parameter $T$ is needed to make the proposed method robust to different datasets whose different numbers of classes.
Table 3: Quantitative Ablation Results of various values of temperature parameter $T$. The experiments are conducted using two typical models including UNet[22] and SegNet[2].
Model $T$ OxfordPetIII CityScapes NyU
@@
SegNet $T=5$ 76.83 91.42 73.86 84.82 19.35 53.39
$T=10$ 76.69 91.40 74.76 85.14 19.49 54.43
$T=20$ 76.69 91.28 76.22 85.80 21.41 55.57
$T=50$ 76.87 91.49 76.21 85.76 21.41 54.29
UNet $T=5$ 78.58 91.42 74.17 84.82 18.23 52.05
$T=10$ 78.74 92.35 74.57 85.34 19.49 54.43
$T=20$ 78.63 92.27 74.85 85.51 21.18 54.22
$T=50$ 78.54 92.29 74.29 85.56 21.32 54.26
Method PAT (Ours) Focal CB CBFocal LDAM BLV
Model UNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
AttUNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NestUNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Image
(b) Label
Figure 5: Visualization of segmentation performance on OxfordPetIII dataset trained by different model architectures including UNet, AttUNet, and NestUNet.
5.2.3 Training performance.
@@
Owing to the demand of taking full advantage of big data which is not only a large scaled number of samples but also high pixel resolution[31], a method that is low cost in both computation and memory usage is essential. To investigate the performance of different methods, we use three metrics including the average training time (seconds/epoch), the average memory acquisition, and the average GPU utilization. We calculate these metrics in each epoch and then take the average value once the training is done.
@@
Fig. 6 suggests that our proposed method (pink circle), which includes the LA and PAT can adapt to a wide range of hardware specifications. While the proposed method acquires roughly 15GB, recent methods (i.e. BLV, LDAM) acquire nearly 17GB and 20GB in the OxfordPetIII and NyU datasets, respectively. In the CityScapes dataset, the proposed method is one of the three lowest GPU-utilized methods, along with the vanilla cross-entropy loss function, which also refers to the lowest time-consumed method.
@@
Figure 6: Performance comparison between baselines and our proposed method in three different scenarios containing OxfordPetIII, CityScapes, and NyU. Performance metrics include Training Time (seconds/epoch), Average Memory Acquisition shown in Gigabyte (GB) units, and the GPU Utilization proportion (%). Note that all experiments are conducted using SegNet[2] model architecture.
5.3 Limitations
Segmentation datasets often exhibit two main types of imbalance including class imbalance (some classes appear more common than others) and size imbalance (some objects account for more proportion than others inside one mask sample)[28]. Through deep empirical analyses, our proposed method can tackle the issues of detecting long-tailed rare objects as well as keeping good performance on well-classified and large proportion categories. However, several key areas still require improvement.
First, although PAT only accounts for a smaller amount of memory, the amount of memory acquired by PAT is still much larger than the Focal and CE. The GPU utilization needed for PAT is not an exception, which is also much higher than Focal and CE, which refers to Fig. 6. This high computation and memory is owing to the large number of calculations needed for the exponential function, though it is a smooth and differentiated function along with its ability to put a high scalarization on low confidence classes.
Second, sharing a common challenge with other semantic segmentation techniques, our proposed method (PAT) is notably susceptible to domain shift. This means that the model’s output predictions (logits) differ significantly across different data domains, making it difficult to determine the optimal scaling loss coefficient. Consequently, applying the model trained on one domain (source) to another domain (target) can result in substantial performance degradation. Future work could address this issue by incorporating domain generalization techniques, such as those described in [23, 32, 14], into the semantic segmentation framework.

### v4

@@ 5 Experimental Evaluations and Discussion
Table 3: Overall Performance of 8 baselines (i.e. Vanilla Softmax, Focal, Class Balance Loss, Class Balance Focal) and proposed method among three different scenarios including OxfordPetIII, CityScape, and NyU datasets. In detail, the bold number indicates the best performance, while $\uparrow$ and $\downarrow$ show ”higher is better” and ”lower is better”, respectively. Note that all experiments shown in this table are trained using SegNet[32] architecture.
@@
CE 76.14 91.21 0.23 73.83 85.31 0.66 18.05 53.50 1.80
Focal ($\gamma=2$) 75.76 91.17 0.30 74.02 85.44 0.56 15.14 51.07 1.75
CB 76.60 90.90 0.20 72.26 81.33 0.69 18.56 52.17 1.94
CBFocal ($\gamma=2$) 76.02 90.54 0.26 71.17 80.70 0.72 17.31 50.46 1.76
BMS 13.22 25.45 1.28 8.15 11.80 3.08 12.27 22.84 2.40
LDAM ($\mu=0.5$, $s=20$) 75.43 90.97 0.78 74.80 85.20 2.27 19.59 52.59 2.30
BLV (Gaussian, $\sigma=0.5$) 76.24 91.22 0.23 74.21 85.37 0.53 18.37 52.62 1.90
@@
76.22 $\uparrow$ 2.2
85.80 $\uparrow$ 0.36
0.51 $\downarrow$ 0.02
21.41 $\uparrow$ 2.85
55.57 $\uparrow$ 2.07
1.36 $\downarrow$ 0.39
We conducted experiments on three popular datasets: OxfordPet[33], CityScapes[34], and NyU[35] whose frequency of classes is considerably sample-wise imbalanced (refers to Section 5.1). The training, validating, and testing ratios are 0.8, 0.1, and 0.1, respectively. OxfordPetIII[33] Dataset contains 37 dog/cat categories ($\approx$200 images/category). Mask ground truth has 3 classes: background, boundary, and main body. Images are original $640\times 40$, resized to $256\times 256$. The biggest obstacle is segmenting the boundary of the animal which is very small and not easily distinguishable. CityScapes[34] dataset contains roughly 5000 images with a high resolution of $1024\times 512$ pixels. This dataset faced a big issue in a high number of imbalanced class imbalances (refers to Fig. 1). The NyU[35] dataset contains roughly 1000 samples whose size of $340\times 256$ pixels. As in the CityScapes dataset, NyU also faces a big challenge in segmenting long-tailed rare objects.
We compare PAT with Focal[15], Class Balance Loss (CB)[16], the combination of CB and Focal (CBFocal)[16], Balance Meta Softmax (BMS)[18], Label Distribution Aware Margin Loss (LDAM)[7], and Balance Logits Variation (BLV)[4] evaluated by mean Intersection over Union (mIoU %), pixel accuracy (Pix Acc %), and Dice Error (Dice Err). The number of rounds of all experiments is fixed to 30000 rounds.
We tuned the hyperparameter of each loss function to find the best case. Specifically, 1) Focal and the combination of Class Balance and Focal are trained with different values of $\gamma\in\{0.5,1,2,3,4,5\}$[15, 16]. 2) In the LDAM, we set the parameter $\mu$ of $0.5$ as the default setting in[17] and trained with different scale $s\in\{10,20,30,50\}$. 3) For the BLV, we applied types of distribution: Gaussian, Uniform, and Xaviver along with different standard deviation $\sigma\in\{0.5,1,2\}$. 4) Hyper-parameter tuning is conducted on PAT with values of $T\in\{5,10,20,50\}$ (refer to Section 5.2.2).
5.1 Comparisons to state-of-the-arts (SOTA)
The metric-based and visual evaluations are shown in Tab. 3 and Fig. 4. In the second column of Fig. 4 (Munich domain), the model trained by PAT can segment fully the road, sky, and most of the building. Visualization performance in Berlin and Leverkusen (refer to Fig. 4), are two typical examples. Other SOTAs tend to misclassify the sky and the road, even though the sky and road owned a considerable proportion.
@@ 5.1 Comparisons to state-of-the-arts (SOTA)
Figure 4: Segmentation visualization of models trained by the PAT and other baselines on the CityScapes dataset.
@@ 5.2 Ablation studies
To guarantee the PAT’s adaptability to different model designs [7, 29, 13, 8, 9, 10], we experiment with different model structures: UNet[36], Attention UNet[37] (AttUNet), Nested UNet[38] (UNet++), DeepLabV3[30] (DLV3), and DeepLabV3+[31] (DLV3+) (refer to Tab. 4). Tab. 4 shows PAT’s superiority on CityScapes, the largest dataset. UNet++ with PAT achieves mIoU and pix acc above 75% and 85%, surpassing BLV and LDAM. Class-Balance (CB) excels on OxfordPetIII (3 classes) but struggles with larger datasets.
Table 4: Comparisons between baselines and PAT in different model architecture designs: UNet, Attention UNet, Nested UNet, DLV3, and DLV3+ whose outperforming cases are indicated by green, blue, yellow, red, and purple colors.
Method Model CityScapes NyU
@@
Focal ($\gamma=2$) UNet 73.53 83.69 15.42 51.03
AttUnet 73.93 83.10 16.46 51.32
UNet++ 74.21 83.52 16.72 51.41
DLV3 77.91 92.78 22.01 57.64
DLV3+ 78.18 93.06 22.35 57.92
CB UNet 78.71 83.19 20.18 54.91
AttUnet 74.47 83.76 19.07 54.66
UNet++ 74.85 83.66 19.17 53.23
DLV3 77.43 92.52 21.86 57.28
DLV3+ 77.77 93.12 22.60 57.96
LDAM ($\mu=0.5$, $s=20$) UNet 73.42 83.81 18.04 54.33
AttUnet 73.73 83.46 19.81 55.22
UNet++ 73.55 84.64 19.04 52.99
DLV3 77.87 92.79 22.28 57.85
DLV3+ 78.40 93.55 23.04 58.63
BLV (Gaussian, $\sigma=0.5$) UNet 74.72 84.74 18.32 54.67
AttUnet 74.45 84.56 19.45 54.72
UNet++ 74.93 84.86 19.81 54.93
DLV3 78.04 93.04 22.40 57.99
DLV3+ 78.42 93.59 23.16 58.71
PAT (Ours) ($T=20$) UNet 74.85 85.51 21.18 54.22
AttUnet 74.57 85.56 21.44 54.41
UNet++ 75.24 85.80 20.66 54.86
DLV3 78.48 93.10 22.66 58.41
DLV3+ 78.63 94.01 23.72 59.09
@@
We perform various experiments of PAT with different temperatures $T\in\{5,10,20,50\}$ (refers to Tab. 5). This ablation test analyzes how temperature parameter $T$ affects the performance of the segmentation model. To make a fair comparison, we conduct all experiments with three related datasets as mentioned in Section 5 with three different types of model architecture including SegNet, UNet, and DLV3+.
Table 5: Quantitative ablation results of various temperatures $T$.
Model $T$ CityScapes NyU
@@
UNet $5$ 74.17 84.82 18.23 52.05
$10$ 74.57 85.34 19.49 54.13
$20$ 74.85 85.51 21.18 54.22
$50$ 74.29 85.56 21.32 54.26
DLV3+ $5$ 78.42 93.66 23.58 58.9
$10$ 78.61 93.76 23.65 59.04
$20$ 78.63 94.01 23.72 59.09
$50$ 79.02 94.45 23.32 59.59
5.2.3 Class-wise performance evaluation
We investigate the class-wise model performance based mIoU metric on CityScapes Dataset using DeepLabV3+ model architecture (refers to Tab. 6). Tab. 6 suggests that PAT can improve the model performance on both head and tail classes.
Table 6: Class-wise experimental evaluation on CityScapes[34] Dataset using DLV3+[31] based mIoU metric. Note that bold and underlined number indicates the highest and second-highest performance cases.
Method
Road
S.Walk
Build.
Wall
Fence
Pole
Light
Sign
Veget.
Terrain
Sky
Person
Rider
Car
Truck
Bus
Train
Motor
Bike
Void
mIoU
CE 99.11 78.60 92.97 63.32 59.07 61.34 64.07 73.82 94.10 52.56 95.60 78.40 56.13 94.89 84.84 85.69 82.17 70.44 68.16 97.79 77.73
Focal 99.48 78.34 92.89 63.62 58.91 61.09 64.45 73.78 94.44 53.04 96.09 78.51 56.55 94.66 85.32 85.52 81.90 70.21 68.34 97.71 78.18
CB 99.33 78.67 92.95 63.58 59.24 61.27 64.02 73.59 94.05 52.70 95.14 78.45 55.96 94.79 85.03 85.55 82.25 70.71 68.97 97.69 77.77
CBFocal 99.37 78.78 93.22 63.28 59.29 61.27 64.54 74.02 94.47 52.55 96.06 78.80 56.40 95.03 85.3 85.84 82.48 70.62 68.90 98.12 78.21
LDAM 99.47 79.60 93.09 63.54 59.73 61.19 65.42 74.40 94.22 52.28 96.76 78.93 56.51 95.17 85.68 85.98 82.54 71.30 68.40 97.84 78.40
BLV 99.12 79.97 93.44 63.12 59.49 61.86 65.52 74.69 95.03 53.13 96.49 79.95 57.36 95.31 86.39 86.35 83.57 71.14 68.83 98.71 78.42
PAT 99.76 79.92 93.89 64.23 60.11 61.11 65.28 75.63 94.95 54.38 95.64 79.19 57.60 95.80 86.52 86.37 84.32 70.23 68.69 99.15 78.63
5.2.4 Training utilization.
@@ 5.2.3 Class-wise performance evaluation
Owing to the demand of taking full advantage of big data which is not only a large scaled number of samples but also high pixel resolution[14, 30, 31], a method that is low cost in both computation and memory usage is essential. To investigate the performance of different methods, we use three metrics including the average training time (seconds/epoch), the average memory acquisition, and the average GPU utilization. We calculate these metrics in each epoch and then take the average value once the training is done.
@@
Fig. 5 suggests that PAT (pink circle), which includes the LA and PAT can adapt to a wide range of hardware specifications. While the proposed method acquires roughly 15GB, recent methods (i.e. BLV, LDAM) acquire nearly 17GB and 20GB in the OxfordPetIII and NyU datasets, respectively. In the CityScapes dataset, the proposed method is one of the three lowest GPU-utilized methods, along with the vanilla cross-entropy loss function, which also refers to the lowest time-consumed method.
@@
Figure 5: Performance comparison between baselines and our proposed method in three different scenarios containing OxfordPetIII, CityScapes, and NyU. Performance metrics include Training Time (seconds/epoch), Average Memory Acquisition shown in Gigabyte (GB) units, and the GPU Utilization proportion (%).